# Zhiwei Zuo Zhuo Tang Kenli Li Anwitaman Datta

### v2

@@ Abstract
Machine unlearning techniques, which involve retracting data records and reducing influence of said data on trained models, help with the user privacy protection objective but incur significant computational costs. Weight perturbation-based unlearning is a general approach, but it typically involves globally modifying the parameters. We propose fine-grained Top-K and Random-k parameters perturbed inexact machine unlearning strategies that address the privacy needs while keeping the computational costs tractable.
In order to demonstrate the efficacy of our strategies we also tackle the challenge of evaluating the effectiveness of machine unlearning by considering the model’s generalization performance across both unlearning and remaining data. To better assess the unlearning effect and model generalization, we propose novel metrics, namely, the forgetting rate and memory retention rate. However, for inexact machine unlearning, current metrics are inadequate in quantifying the degree of forgetting that occurs after unlearning strategies are applied. To address this, we introduce SPD-GAN, which subtly perturbs the distribution of data targeted for unlearning. Then, we evaluate the degree of unlearning by measuring the performance difference of the models on the perturbed unlearning data before and after the unlearning process. By implementing these innovative techniques and metrics, we achieve computationally efficacious privacy protection in machine learning applications without significant sacrifice of model performance. Furthermore, this approach provides a novel method for evaluating the degree of unlearning.
keywords:
Machine Unlearning, Perturbation, User Privacy, Forgetting
[…]
1 Introduction
Users may want the removal of personal data in possession of an organization. In several jurisdictions, this has compliance implications under regulations such as the General Data Protection Regulation (GDPR) in the European Union goddard2017eu and the California Consumer Privacy Act in the United States goldman2020introduction . Machine learning models trained using such data retain their influence, thus risking privacy compromises cheng2023privacy ; wu2023privacy . For instance, in a recommendation system using collaborative filtering schafer2007collaborative ; isinkaye2015recommendation , if the model does not eliminate the influence of removed data (which we call ‘unlearning data’), it may continue to make recommendations based on user similarity.
@@ 1 Introduction
Exact unlearning thudi2022necessity ; bourtoule2021machine aim to completely eliminate the influence of unlearning data on a model. Some researches try to evaluate the impact of data on models, the Influence function offers insights but faces limitations in deep learning contexts koh2017understanding ; basu2020influence ; li2023selective . In contrast, the Single-Step Sample Erasure (SSSE) peste2021ssse utilizes the Fisher Information Matrix to sidestep computational hurdles. In contrast, inexact machine unlearning DBLP:journals/corr/abs-2201-06640 ; suriyakumar2022algorithms ; neel2021descent aims to reduce time and resources requirements by selectively eliminating the data’s influence on the model. Current inexact machine unlearning methods based on model weight perturbation typically involve adding random Gaussian noise to all model parameters. We want to conduct unlearning using a more fine-gained perturbation. To that end, we design two strategies: Random-k and Top-K. By selectively perturbing a small subset of parameters, we aim to achieve the desired effect of machine unlearning efficiently while also reducing the impact on the overall model performance.
In the area of machine unlearning, model indistinguishability golatkar2020eternal ; kurmanji2023towards is crucial to consider. When different algorithms learn from the same data set, they may acquire very similar knowledge and features. Consequently, when faced with independent and identically distributed (i.i.d) datasets, even if an unlearning approach has eliminated the influence of unlearning data, the model may still show approximate accuracy on the unlearning data to the remaining data due to the model’s generalization property nakkiran2020distributional . This makes it challenging to evaluate the effectiveness of the unlearning process.
Consequently, to measure the effectiveness of unlearning, metrics can be divided into two aspects. The first is general performance evaluation, which includes assessing the accuracy marchant2022hard of unlearning model predictions, unlearn time corresponding to the unlearning request bourtoule2021machine ; becker2018organizational , relearn time for reaching the accuracy of source model tarun2021fast , membership inference attack (MIA) shokri2017membership ; hu2022membership . The second focuses on model indistinguishability, completeness, activation distance wu2020deltagrad , Jensen–Shannon (JS)-Divergence chundawat2023zero , and other metrics that quantify the differences between the unlearning and retraining models.
In the fine-grained inexact machine unlearning strategies we employ, we consider both the impact of model unlearning and the generalization performance of the model. A metric called forgetting rate is designed to assess the effect of model unlearning, while memory retention rate and similarity are employed to evaluate the model’s indistinguishability.
@@
However, the existing metrics are unable to reveal the extent of unlearning. Suppose splitting training data into two parts: unlearning data $D_{UL}$ and remaining data $D_{RE}$ $\left(D=D_{UL}\cup D_{RE},len(D_{UL})\ll len(D_{RE})\right)$. As shown in Fig.1, while the accuracy may vary across $D_{UL}$ and $D_{RE}$ for each model, it consistently maintains a high accuracy for $D_{UL}$. Consequently, it becomes challenging to quantify the extent to which the model eliminates the influence of $D_{UL}$. Since the performance of unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$ is relatively small, in order to amplify the performance difference between the unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$, we propose a SPD-GAN (Slightly Perturbed Distribution-Generative Adversarial Network) to break the i.i.d attribute of the unlearning data. For the unlearning data $D_{UL}$, we introduce perturbations to it’s distribution, thus we gain perturbed unlearning data $D_{p}$. We limit the perturbations to be very slight so that the accuracy of source model $M$ on $D_{p}$ is close to the accuracy of $M$ on $D_{UL}$. Suppose $P(M,D)$ represents the performance of model $M$ on data $D$, seeing in the Equation.1, if the perturbed unlearning data $D_{p}$ exhibits significant performance difference between source model and unlearning model, we can then obtain the unlearning degree of each unlearning strategies.
$\centering P(M,D_{p})\gg P(M_{UL},D_{p})\quad s.t.\quad P(M,D_{p})\approx P(M,D_{UL})\@add@centering$ (1)
@@
1. We adopt two fine-grained inexact machine unlearning strategies. We propose strategies of Random-k and Top-K to quickly achieve the unlearning effect with minimal model parameters perturbed while ensuring its generalization properties (Section 3.2).
2. We design measurement metrics forgetting rate, memory retention rate and similarity to assess unlearning effectiveness and the unlearning model’s generalization properties. We carry out a theoretical analysis in Section 3.4 of the acceleration of inexact machine unlearning strategies, and represent it as acceleration ratio in our experiments (as discussed in Section 4.3.3).
@@
3. We design a novel unlearning degree evaluation method by proposing a SPD-GAN to slightly perturb the unlearning data, using the performance difference of the perturbed unlearning data on the source model and unlearning model to approximate the degree of unlearning (Section 3.3). Experimental results show that our proposed Top-K can achieve the deepest degree of unlearning in comparison to other inexact unlearning strategies (as discussed in Section 4.3.5).
2 Related Works
2.1 User Privacy
Several important existing methods aim to protect user privacy. Traditional approaches focus on decoupling and generalization to eliminate user identifiers or use anonymization techniques bayardo2005data to generalize or compress user information. However, these methods may not be effective when prior knowledge is available.
Differential privacy (DP) dwork2006calibrating is a strict and provable privacy protection technique that adds noise in the stochastic gradient descent(SGD) process. Other approaches, such as dPA kocher2011introduction , add perturbation to the objective function, while PATE papernot2018scalable uses knowledge aggregation and transmission. However, the process of adding noise in DP is irreversible, making it unsuitable for machine unlearning tasks since the protection cannot be removed once data is protected by noise.
@@ 2.1 User Privacy
Federated learning (FL) is another method to protect user privacy, where training data is distributed among multiple devices and servers, and the model is trained in a decentralized manner without centralizing the data wei2020federated . FL methods based on differential privacy have been proposed to protect user privacy, with the principal focus on attacks xiao2022sca and defenses xiao2022sbpa in FL systems.
2.2 Machine Unlearning
@@
Several approaches have emerged to address the limitations of retraining models from scratch for machine unlearning, since it is generally impractical. The typical approach is SISA (Sharded, Isolated, Sliced, and Aggregated) training, which splits data into shards and trains them in isolation for speed and effectiveness when unlearning. liu2022continual combines continual learning with private learning based on the SISA idea, but SISA-based methods often consume a lot of storage.
@@
Other approaches focus on evaluating the impact of specific data points or sub-datasets on trained models. Influence function koh2017understanding approximates the parameter change when data is up-weighted by a small value, but its effectiveness in deep learning is limited by non-convex loss functions and Taylor’s approximation errors basu2020influence . Peste et al. propose a Single-Step Sample Erasure (SSSE) using the Fisher Information Matrix (FIM) to avoid expensive computation and Hessian inversion peste2021ssse . Some works use influence functions in Graph Neural Networks (GNNs) for unlearning, but these models are relatively shallow, and Hessian inversion is more accurate, allowing for exact unlearning.
Many model-related methods have been developed for machine unlearning. Ganhor et al. ganhor2022unlearning incorporate adversarial training into multVAE liang2018variational to remove implicit information of protected attributes. Liu et al. liu2022right study unlearning in federated learning and propose a fast retraining method to erase the impact of data samples from trained models. Some works discuss unlearning strategies in Bayesian methods rawat2022challenges , while others focus on random forests liu2021revfrf .
2.3 Adversarial Learning
Adversarial learning is an adversarial-based learning method that aims to enhance the robustness of machine learning algorithms against attacks. However, it may also lead to user privacy leakage, which can have severe consequences in areas such as biometric authentication.
The core concept of adversarial learning involves introducing minor perturbations to input data in order to cause incorrect classification. Szegedy et al. szegedy2013intriguing reveal the vulnerability of neural networks to adversarial examples and introduce the L-BFGS attack method. Due to the high time consumption associated with L-BFGS, Goodfellow et al. goodfellow2014explaining propose the fast gradient sign method (FGSM) which optimizes $L_{\infty}$distance. This method is later refined by Carlini and Wagner into the $C\&W$ method carlini2017towards . Additionally, Papernot et al. papernot2016limitations propose the Jacobian-based saliency map attack (JSMA), which optimizes under $L_{0}$ distance.
@@ 2.3 Adversarial Learning
The concept of perturbation through the addition of noise is also employed in machine unlearning tasks. neel2021descent utilizes random Gaussian noise perturbations to the source model $M$’s parameters, constrained by ($\epsilon,\delta$), ensuring that the model remains indistinguishable from any point within a small neighborhood of the optimal model.
@@
3 Methodology
3.1 Problem Definition
@@
In this paper, we focus on machine unlearning problem. Let $D$ represent the training data set which contains $n$ samples $\left(x_{1},y_{1}\right),\left(x_{2},y_{2}\right),...,\left(x_{n},y_{n}\right)$. For each sample, $x\in X$ and $y\in Y$. These data are posted by multiple users and collected by some institutions or organization and then labeled. Suppose $D=D_{UL}\cup D_{RE}$ and $D_{UL}\cap D_{RE}=\emptyset$, here $D_{RE}$ represents the remaining data after unlearning data removal. Given a trained model $M$ with parameter $w$ on data set $D$, when conducting machine unlearning, our goal is to gain a target model $M_{UL}$ with parameter $\theta$, which eliminates the influence of $D_{UL}$ and holds a quite good performance on $D_{RE}$.
@@
We assume the posterior distribution of $M$ and $M_{UL}$ are $M\left(D_{UL}\right)\sim P_{1}$ and $M_{UL}\left(D_{UL}\right)\sim P_{2}$. When $P_{1}$ and $P_{2}$ are quite similar, we claim machine unlearning is conducted.
3.2 Our Approach

### v3

# Machine Unlearning Through Fine-Grained Model Parameters Perturbation

@@ Abstract
Machine unlearning involves retracting data records and reducing their influence on trained models, aiding user privacy protection, at a significant computational cost potentially. Weight perturbation-based unlearning is common but typically modifies parameters globally. We propose fine-grained Top-K and Random-k parameters perturbed inexact machine unlearning that address the privacy needs while keeping the computational costs tractable.
However, commonly used training data are independent and identically distributed, for inexact machine unlearning, current metrics are inadequate in quantifying forgetting degree that occurs after unlearning. To address this quantification issue, we introduce SPD-GAN, which subtly perturbs data distribution targeted for unlearning. Then, we evaluate unlearning degree by measuring the performance difference of the models on the perturbed unlearning data before and after unlearning. Furthermore, to demonstrate efficacy, we tackle the challenge of evaluating machine unlearning by assessing model generalization across unlearning and remaining data. To better assess the unlearning effect and model generalization, we propose novel metrics, namely, the forgetting rate and memory retention rate. By implementing these innovative techniques and metrics, we achieve computationally efficacious privacy protection in machine learning applications without significant sacrifice of model performance. A by-product of our work is a novel method for evaluating and quantifying unlearning degree.
Index Terms:
Inexact Machine Unlearning, Perturbation, User Privacy, Forgetting, Unlearning Degree Quantification
I Introduction
Users may want the removal of personal data in possession of an organization, typically deleting it randomly based on their specific needs. In several jurisdictions, this has compliance implications under regulations such as the General Data Protection Regulation (GDPR) in the European Union [1] and the California Consumer Privacy Act in the United States [2]. Machine learning models trained using such data retain their influence, thus risking privacy compromises [3]. For instance, in a recommendation system using collaborative filtering [4], if the model does not eliminate the influence of removed data (which we call ‘unlearning data’), it may continue to make recommendations based on user similarity.
@@ I Introduction
Exact machine unlearning methods [5, 6] are designed to completely eliminate the influence of unlearning data on a model. Although some methods explore the impact of data on models, the Influence Function provides insights into such impacts, but it has limitations. The non-convexity of the loss functions used in deep learning models leads to significant approximation errors, and the computational burden of calculating the inverse of the Hessian matrix in the Influence Matrix limits its application in deep learning environments [7, 8]. In contrast, the Single-Step Sample Erasure (SSSE) [9] utilizes the Fisher Information Matrix to sidestep computational hurdles, achieving a more precise approximation.
Differently, inexact machine unlearning [10, 11] aims to reduce time and resources requirements by selectively eliminating the data’s influence on the model. Current inexact machine unlearning methods based on model weight perturbation typically involve adding random Gaussian noise to all model parameters. We propose to achieve inexact unlearning using a more fine-gained perturbation. To that end, we design two inexact machine unlearning strategies: Random-k and Top-K. By selectively perturbing a small subset of parameters, we aim to achieve the desired effect of machine unlearning efficiently while also reducing the impact on the overall model performance.
Even if we conduct machine unlearning successfully, evaluating unlearning effectiveness remains a challenge, especially in terms of how to quantify the degree of unlearning. This is because when different algorithms learn from the same data set, they may acquire very similar knowledge and features. Consequently, when faced with independent and identically distributed (i.i.d) datasets, even if an unlearning approach has eliminated the influence of unlearning data, the model may still show approximate accuracy on the unlearning data to the remaining data due to the model’s generalization property [12]. This makes it challenging to evaluate the effectiveness of the unlearning process.
In practice, to measure the effectiveness of unlearning, current metrics can be divided into two aspects. The first is general performance evaluation, which includes assessing the accuracy [13] of unlearning model predictions, unlearn time corresponding to the unlearning request [6], relearn time for reaching the accuracy of source model [14], membership inference attack (MIA) [15, 16]. The second focuses on model indistinguishability [17, 18], completeness, activation distance [19], Jensen–Shannon (JS)-Divergence [20], and other metrics that quantify the differences between the unlearning and retraining models.
@@
However, the existing metrics are unable to reveal the extent of unlearning. Suppose splitting training data into two parts: unlearning data $D_{UL}$ and remaining data $D_{RE}$ $\left(D=D_{UL}\cup D_{RE},len(D_{UL})\ll len(D_{RE})\right)$. As shown in Fig.1, which shows an early glimpse of some results achieved by one of our own unlearning techniques), while the accuracy may vary across $D_{UL}$ and $D_{RE}$ for each model, it consistently maintains a high accuracy for $D_{UL}$. Consequently, it becomes challenging to quantify the extent to which the model eliminates the influence of $D_{UL}$. Since the performance difference of unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$ is relatively small. We attribute this to two reasons, first, a good generalization capability of the model, second, the characteristic that the training data followed an independent and identically distributed (i.i.d) pattern.
In order to amplify the performance difference between the unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$, we propose a SPD-GAN (Slightly Perturbed Distribution-Generative Adversarial Network) to break the i.i.d attribute of the unlearning data. For the unlearning data $D_{UL}$, we introduce perturbations to it’s distribution, thus we gain perturbed unlearning data $D_{p}$. We limit the perturbations to be very slight so that the accuracy of source model $M$ on $D_{p}$ is close to the accuracy of $M$ on $D_{UL}$. Suppose $P(M,D)$ represents the performance of model $M$ on data $D$, seeing in the Equation.1, if the perturbed unlearning data $D_{p}$ exhibits significant performance difference between source model and unlearning model, we can then obtain the unlearning degree of each unlearning strategies.
$P(M,D_{p})\gg P(M_{UL},D_{p})$
$s.t.~{}P(M,D_{p})\approx P(M,D_{UL})$ (1)
In addition to measuring the degree of unlearning, in the fine-grained inexact machine unlearning strategies we employ, we also consider both the impact of model unlearning and the generalization performance of the model. A metric called forgetting rate is designed to assess the effect of model unlearning, while memory retention rate and similarity are employed to evaluate the model’s indistinguishability.
@@
  • We adopt two fine-grained inexact machine unlearning strategies of Random-k and Top-K to quickly achieve the unlearning effect with minimal model parameters perturbed instead of perturbed globally while ensuring its generalization properties (Section III-B).
@@
  • We design a novel unlearning degree evaluation method by proposing a SPD-GAN to slightly perturb the unlearning data, using the performance difference of the perturbed unlearning data on the source model and unlearning model to approximate the degree of unlearning (Section III-C). Experimental results show that our proposed Top-K can achieve the deepest degree of unlearning in comparison to other inexact unlearning strategies (as discussed in Section IV-C5).
  • We design measurement metrics forgetting rate, memory retention rate and similarity to assess unlearning effectiveness and the unlearning model’s generalization properties. We carry out a theoretical analysis in Section III-D of the acceleration of inexact machine unlearning strategies, and represent it as acceleration ratio in our experiments (as discussed in Section IV-C3).
In the subsequent sections, we first explore the related work in Section II, then describe the methodology in Section III, report the findings from our experiments in Section IV, and finally, draw conclusions and discuss future research directions in Section V.
II Related Works
II-A User Privacy
For data-driven decision-making systems, the data used is primarily collected from the internet or exposed by users when they interact with such systems, much of which is highly sensitive and confidential. There are many methods trying to serve the function of privacy protection in various scenarios. Traditional approaches achieve decoupling by separating parts of the data that are directly related to the user, perform generalization by erasing or replacing specific details in the data or protect privacy using other anonymization techniques such as data desensitization and data perturbation [21]. However, these methods may not be effective when prior knowledge is unavailable.
Differential privacy (DP) [22] provides a strict definition of privacy, making it the de facto standard across various data types in the field of privacy protection. Differential privacy (DP) operates by adding well-calibrated noise to individual data points or database queries, ensuring that the results of data analysis do not significantly change due to the presence or absence of any single data point. Local differential privacy (LDP) [23] perturbs data before it leaves a user’s local device, ensuring that only the data owner retains access to the original data. However, LDP can introduce excessive noise into the dataset. Other approaches, such as dPA [24] add perturbation to the objective function, while PATE [25] uses knowledge aggregation and transmission. However, the process of adding noise in DP is irreversible, making it unsuitable for machine unlearning tasks since the protection cannot be removed once data is protected by noise, and more crucially in our context, it also is not suitable for removing influence of data from a model that has already been trained with a given data.
@@ II-A User Privacy
Federated learning (FL) is a decenterlized approach to protect user privacy, where training data is distributed among multiple devices and servers, and the model is trained in a decentralized manner without centralizing the data [26]. FL methods based on differential privacy have been proposed to protect user privacy, with the principal focus on attacks and defenses [27] in FL systems. Furthermore, recent advancements in FL incorporate various cryptographic techniques, such as secure multi-party computation and homomorphic encryption, to enhance the security and privacy of the training process. This ongoing research strengthens the privacy-preserving capabilities of FL.
Other decentralized privacy protection technologies, like blockchain [28], distribute data across many different nodes in the network to reduce the risk of data tampering or unauthorized access, while also using encryption technologies to protect the transmission and access of data.
II-B Machine Unlearning
@@
Several approaches have emerged to address the limitations of retraining models from scratch for machine unlearning, since it is generally impractical. The most typical approach is SISA (Sharded, Isolated, Sliced, and Aggregated) training [6], which splits data into shards and trains them in isolation for speed and effectiveness when unlearning. Liu et al. [29] combine continual learning with private learning based on the SISA idea, but SISA-based methods often consume a lot of storage. Following the combination of machine unlearing and continual learning, Zuo et al. [30] propose a embedding-based framework, which modifies the data rather than the model itself to achieve unlearning goal by using vector databases.
@@
Other approaches focus on evaluating the impact of specific data points or sub-datasets on trained models. Influence function [31] approximates the parameter change when data is up-weighted by a small value, but its effectiveness in deep learning is limited by non-convex loss functions and Taylor’s approximation errors [7]. Peste et al. propose a Single-Step Sample Erasure (SSSE) using the Fisher Information Matrix (FIM) to avoid expensive computation and Hessian inversion [9]. Some works use influence functions in Graph Neural Networks (GNNs) for unlearning, but these models are relatively shallow, and Hessian inversion is more accurate, allowing for exact unlearning.
In addition to combining with continual learning, machine unlearning is also integrated with other approaches, such as federated learning, forming a research area known as federated unlearning [32] [33]. In federated unlearning, considerations are made for clients leaving the federation or for unlearning data corresponding to clients. For instance, Liu et al. [32] propose FedEraser, which unlearns a client’s data by reconstructing an unlearned global model instead of retraining from scratch. Wu et al. [33] utilize knowledge distillation techniques to recover the contribution of client models to achieve federated unlearning.
While the strategy for unlearning is crucial, the verification of unlearning is also a critical issue. Membership Inference Attack [34, 35] is a common verification method. Utilizing the characteristics of backdoor attacks, users embed a specific backdoor trigger into the data, and then use the backdoor attack to test whether the model has been trained on this data or has successfully eliminated the influence of this data.
II-C Adversarial Learning
Even though neural networks are widely applied, their robustness is hard to ensure. An adversarial sample is the data generated by adding slight perturbation $\delta x$ to raw data $x$, which can cause the neural network to misclassify yet human can still make the correct classification. This leads to potentially fatal dangers in critical areas such as self-driving or biometric authentication, where robust models are essential.
For perturbation generation, Szegedy et al. [36] reveal the vulnerability of neural networks to adversarial examples and introduce the L-BFGS (Limited Memory Broyden-Fletcher-Goldfarb-Shanno) attack method. Due to the high time consumption associated with L-BFGS, Goodfellow et al. [37] propose the fast gradient sign method (FGSM) which optimizes $L_{\infty}$ distance. This method is furthermore refined by Carlini and Wagner into the $C\&W$ method [38]. Additionally, Papernot et al. [39] propose the Jacobian-based saliency map attack (JSMA), which optimizes under $L_{0}$ distance. DeepFool [40] looks for the minimal change necessary to deceive the model into misclassifying an input under the specific condition that the perturbation vector is orthogonal to the hyperplane representing the classifier’s decision boundary. Generative Adversarial Network (GAN) [41] is also a promising approach to generate adversarial samples. Each of these methods demonstrates different aspects of the neural networks’ vulnerabilities and their respective optimizations highlight various approaches to improving attack efficiency.
Moreover, these techniques have laid the groundwork for numerous subsequent studies in adversarial machine learning, illustrating the continuous evolution of this field. In other words, leveraging adversarial training enhances the robustness of the model and better identify and defend against malicious attacks. Smoothed ViTs [42] are trained using only one columns of original data while the remaining are black. Pyramid Adversarial Training [43] generates adversarial samples and perform a matched Dropout technique and stochastic depth regularization to enhance the robustness.
@@ II-C Adversarial Learning
The concept of perturbation through the addition of noise is also employed in machine unlearning tasks. [44] utilizes random Gaussian noise perturbations to the source model $M$’s parameters, constrained by ($\epsilon,\delta$), ensuring that the model remains indistinguishable from any point within a small neighborhood of the optimal model.
@@
III Methodology
III-A Problem Definition
@@
In this paper, we focus on machine unlearning problem. Let $D$ represent the training data set which contains $n$ samples $\left(x_{1},y_{1}\right),\left(x_{2},y_{2}\right),...,\left(x_{n},y_{n}\right)$. For each sample, $x\in X$ and $y\in Y$. These data are posted by multiple users and collected by some institutions or organization and then labeled. Suppose $D=D_{UL}\cup D_{RE}$ and $D_{UL}\cap D_{RE}=\emptyset$, here $D_{UL}$ represents the unlearning data, which is absolutely randomly chosen by users to be unlearned and $D_{RE}$ represents the remaining data after unlearning data removal. Given a trained model $M$ with parameter $w$ on data set $D$, when conducting machine unlearning, our goal is to gain a target model $M_{UL}$ with parameter $\theta$ while the number of perturbed parameters should be as few as possible, and this $M_{UL}$ eliminates the influence of $D_{UL}$ and holds a quite good performance on $D_{RE}$. Suppose $w^{\prime}$ is the parameter after fine-grained perturbation. And as previously mentioned, we assume $P(M,D)$ as the performance of a model $M$ on data $D$, thus, the effect that our fine-grained perturbation machine unlearning aims to is :
$P(M_{\theta},D_{UL})\ll P(M_{\theta},D_{RE})$
$P(M_{\theta},D_{RE})\approx P(M_{w},D_{RE})$
$minize\|w^{\prime}-w\|_{0}$ (3)
where $\|w^{\prime}-w\|_{0}$ is $L_{0}$ norm, representing the number of perturbed parameters.
@@
Our fine-grained perturbation machine unlearning includes two stages: perturbation and fine-tuning. And the similarity of the output distributions of the model $M$ and $M_{UL}$ guides whether the fine-tuning should be terminated. We assume the posterior distribution of $M$ and $M_{UL}$ are $M\left(D_{UL}\right)\sim P_{1}$ and $M_{UL}\left(D_{UL}\right)\sim P_{2}$. When $P_{1}$ and $P_{2}$ are quite similar, we claim machine unlearning is conducted.
III-B Our Unlearning Approachs